what is the OH- activity (concentration),[OH-], of waterwith a pH of 4? Express your answer in moles/L and mg/L.
Answer #1

pH = -log[aH+]

pOH = 14 - pH

pOH = 14 -4 = 10

log [aOH-] = -10

[aOH-] = 10^-10 mol / L

[aOH-] = 10^-10 * 17 g/mol

[aOH-] = 1.7 * 10^-6 mg / L